# SECURE: Benchmarking Generative Large Language Models for Cybersecurity Advisory

### v1

@@ Abstract.
Large Language Models (LLMs) have demonstrated potential in cybersecurity applications but have also caused lower confidence due to problems like hallucinations and a lack of truthfulness. Existing benchmarks provide general evaluations but do not sufficiently address the practical and applied aspects of LLM performance in cybersecurity-specific tasks. To address this gap, we introduce the SECURE (Security Extraction, Understanding & Reasoning Evaluation), a benchmark designed to assess LLMs performance in realistic cybersecurity scenarios. SECURE includes six datasets focussed on the Industrial Control System sector to evaluate knowledge extraction, understanding, and reasoning based on industry-standard sources. Our study evaluates seven state-of-the-art models on these tasks, providing insights into their strengths and weaknesses in cybersecurity contexts, and offer recommendations for improving LLMs reliability as cyber advisory tools.
Large Language Models, Security, Cyber Advisory, Dataset
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Annual Computer Security Applications Conference (ACSAC); December 09–13, 2024; Waikiki, HIisbn: 978-1-4503-XXXX-X/24/12ccs: Security and privacy Usability in security and privacyccs: Human-centered computing Heuristic evaluations
1. Introduction
Figure 1. Conversation between a user and a ChatGPT-3.5 on ICS security questions showing incorrect responses from the language model.
@@ 1. Introduction
Recent breakthroughs in large language models (LLM) like OpenAI’s ChatGPT (gpt, 2024b) have opened up applications of generative models in many domains, including security (Zaboli et al., 2023). These models, trained on vast datasets, can generate, understand, and reason across a multitude of domains (Bubeck et al., 2023). Users can interact with these models in a conversational style through a simple web interface and obtain answers to their questions in a short time. Despite demonstrating high potential, LLMs are plagued by issues such as hallucinations (Weidinger et al., 2021) and truthfulness (Madaan and Yazdanbakhsh, 2022). Hence, a standard evaluation benchmark is crucial to evaluate the reliability of such models. Several benchmarks like GLUE (Wang et al., 2018), MMLU (Hendrycks et al., 2020), Helm (Liang et al., 2022) and KOLA (Yu et al., 2023), provide standard datasets and tasks to evaluate general-purpose understanding and capabilities of LLMs. GLUE assesses LLMs’ performance in understanding language, while MMLU and HELM offer a holistic evaluation across various domains like mathematics, history, computer science, and law. KOLA focuses on tasks designed to measure the cognitive abilities of LLMs.
@@
Despite these advancements, there remains a significant gap in the evaluation of LLMs specifically tailored for security industries such as information security, network security, and critical infrastructure protection. Traditional benchmarks often fail to capture the practical and applied aspects of cybersecurity, leading to an incomplete assessment of LLM capabilities. These benchmarks typically focus on general language tasks and do not address specific challenges such as recognizing emerging threats, handling specialized terminology, or performing tasks like vulnerability assessment and incident response. More practical, domain-specific, and comprehensive evaluations are necessary to understand LLM performance in realistic cybersecurity scenarios (Gennari et al., 2024). In Figure 1, we demonstrate a conversation between ChatGPT-3.5 and a user based on our benchmark dataset (see Section 3). ChatGPT-3.5 model was prompted to act as a security expert, and yet, all three responses, spanning different kinds of tasks, were incorrect, showing the unreliability of these models in cybersecurity. Especially, on the first task, we ask questions about vulnerabilities discovered in 2024. Even though the model is not trained on any such recent data, it confidently makes a decision, despite our instructions to return ’X’ when it is not confident.
@@
To address this gap, we introduce a comprehensive benchmarking framework encompassing real-world cybersecurity scenarios, practical tasks, and applied knowledge assessments. We introduce Security ExtraCtion, Understanding & Reasoning Evaluation (SECURE) for large language models. This benchmark provides a holistic evaluation of LLMs performance in cybersecurity applications, ensuring they meet the high standards required for deployment in critical infrastructure environments, such as Industrial Control Systems (ICS). Figure 2 shows an overview of our proposed benchmark. SECURE incorporates knowledge modeling to design six different datasets (MAET: Mitre Attack Extraction Task, CWET: Common Weakness Extraction Task, KCV: Knowledge test on Common Vulnerabilities, VOOD: Vulnerability Out-of-Distribution task, RERT: Risk Evaluation Reasoning Task, and CPST: CVSS Problem Solving Task) focused on three types of knowledge evaluations: extraction, understanding, and reasoning. These datasets are sourced from standard sources, such as MITRE (mit, 2024c) (mit, 2024a), CVE (Project, 2024a), CWE (Project, 2024b), and Cybersecurity and Infrastructure Security Agency (CISA) (Advisories, 2024).
@@
    Assessment of cybersecurity knowledge: We evaluate the cybersecurity knowledge of general-purpose LLMs and their reliability in assisting security analysts. To achieve this, we develop a series of benchmark tasks specifically for evaluating LLMs in the context of ICS cyber advisory (see Section 3). These tasks are based on comprehensive protocol specifications detailing communication standards and procedures, historical cyber threats and attacks, CVEs with contextual information on exploitability and severity, and detailed remediation strategies, including patch notes and mitigation steps from MITRE.
  2. (2)
    Evaluation of LLMs: We evaluate 7 open and closed source state-of-the-art models (ChatGPT-4 (gpt, 2024b), ChatGPT-3.5 (gpt, 2024a), Llama3-70b (lla, 2024a), Llama3-8b (lla, 2024b), Gemmini-Pro (gemmini), Mistral-7B (Jiang et al., 2023) , Mixtral-8x7b (mix, 2024)) on our benchmark tasks (see Section 4).
  3. (3)
@@
    Insights and Recommendations: Our findings demonstrate that while LLMs demonstrate some capability in cybersecurity tasks, their use as cyber advisory tools requires careful consideration. We provide insightful observations and propose recommendations to enhance their usability (see Section 5).
Figure 2. Overview of the SECURE benchmark.
The remainder of this paper is structured as follows: Section 2 reviews background and related work in the application of LLMs in cybersecurity. Section 3 introduces the benchmark and associated tasks. Section 4 discusses the experiments and results. Section 5 presents our analysis of the results and its implications. Section 6 discusses our limitations and future work, with conclusion of the paper in Section 7.
2. Background and Related Work
Generative models are capable of creating new content, such as text, images, video, and other forms of data, by learning patterns from existing datasets. Such models are trained on extensive datasets, enabling them to generate output that mimics the style and structure of the training data. This has led to applications across diverse fields such as text generation, and image creation (Harshvardhan et al., 2020). While the potential of generative AI is vast, there are challenges such as ensuring content quality, coherence, and reliability. Below, we provide a background on such generative large language models in security and approaches to evaluate them.
@@ 2. Background and Related Work
Generative Large Language Model in Security: Generative Large Language Models (LLMs) are designed to understand and generate human language. These models are trained on vast amounts of textual data and are capable of producing coherent and contextually relevant text. They leverage deep learning techniques, particularly neural networks with transformer architectures (Vaswani et al., 2017), to capture the nuances of language. While earlier BERT-language models like SecureBERT (Aghaei et al., 2022) and CySecBERT (Bayer et al., 2024) were used for language modeling in cybersecurity, the release of GPT (Generative Pre-trained Transformer) models has changed the nature of language models. There are two types of LLMs in the industry. Open Source LLMs like Llama (lla, 2024a) and Mixtral (Jiang et al., 2023) make their models public so they can be fine-tuned for specific downstream tasks. Closed-source LLMs like ChatGPT (gpt, 2024b) and Gemini (Team et al., 2024) allow restricted access through APIs. The most notable language models in security are code-based LLMs such as CodeLlama (Roziere et al., 2023), adapted to analyze and generate secure code. Open-sourced models have been also fine-tuned for several cybersecurity tasks like vulnerability detection (Shestov et al., 2024; Ferrag et al., 2023), program repair (Silva et al., 2023), IT operations (Guo et al., 2023), and security knowledge assistance (Sultana et al., 2023).
@@
Evaluation Benchmark: While there are various benchmarks like GLUE (Wang et al., 2018), MMLU (Hendrycks et al., 2020), Helm (Liang et al., 2022) and KOLA (Yu et al., 2023) for evaluating general-purpose LLMs, comprehensive cybersecurity-specific benchmarks remain limited. Existing approaches to evaluating LLMs in security tend to focus on factual knowledge rather than applied, practical cybersecurity tasks. For example: the CyberMetric dataset (Tihanyi et al., 2024) comprises 10,000 questions sourced from cybersecurity standards, certifications, and publications, created through a collaborative process involving experts and LLMs. CyberBench (Liu et al., [n. d.]) comprises ten datasets from different tasks namely, named-entity recognition, summarization, multiple choice, and classification. SecEval (Li et al., 2023) evaluates cybersecurity knowledge in LLMs with 2000 multiple-choice questions across Software Security, Application Security, System Security, Web Security, Cryptography, Memory Safety, Network Security, and PenTest. SecQA (Liu, 2023) consists of multiple-choice questions based on the ”Computer Systems Security: Planning for Success” textbook to evaluate LLM’s understanding of security principles. NetEval (Miao et al., 2023) evaluates the knowledge of large language models in IT operational tasks within a multilingual context. OpsEval (Liu et al., 2023) also contains multi-choice questions designed for fault root cause analysis, operational script generation, and alert information summarization, to assess LLMs in IT operational tasks. These benchmarks are however limited to evaluating memorization of the LLMs as such questions only measure the information retrieval ability of large language models; and lack an evaluation of the comprehension and problem-solving capabilities of LLMs in cybersecurity.
3. Proposed Benchmark: SECURE
@@ 3.1. Modeling
We emphasize knowledge modeling in designing our evaluation benchmark. Knowledge, which includes both facts and skills, is a core indicator of intelligence (Feigenbaum et al., 1977). Prior research shows that knowledge-intensive tasks can reliably evaluate the capabilities of LLMs (Yu et al., 2023; Petroni et al., 2019). For cyber-advisory LLM, we aim to assess not only its ability to retrieve known facts but also gauge its proficiency in a) using context to answer questions and b) performing reasoning based on given some knowledge source. To create a robust evaluation benchmark, we focus on the following critical abilities of LLMs in handling knowledge-intensive tasks:
@@
Extraction: Knowledge extraction tasks are crucial for measuring the ability of a language model to access its vast knowledge base and accurately recall specific facts (Petroni et al., 2019). As a cybersecurity advisor, an LLM could be asked to retrieve information on various security architectures, incidents, best practices, and historical data on known vulnerabilities. Accurate extraction ensures the LLM can provide reliable and precise information quickly, which is essential for addressing security concerns, preventing breaches, and supporting security professionals in their decision-making processes.
@@
Understanding: While knowledge extraction tasks focus on recalling information, knowledge understanding tasks are designed to assess the cognitive abilities of a model (Collins and Olson, 2014). This involves evaluating the model’s capability to discern the truthfulness of statements and comprehend underlying knowledge within a given context. For instance, an LLM might be tested on its ability to interpret the accuracy of security issues described in a report, thereby demonstrating its grasp of complex cybersecurity concepts and scenarios. Effective understanding ensures that the LLM can accurately interpret and respond to nuanced security challenges.
@@
Reasoning: Knowledge reasoning tasks aim to evaluate the problem-solving capabilities of LLMs (Cao et al., 2022). This is particularly important in the cyber-advisory role of LLMs, as they need to assist security professionals in reading, analyzing, and summarizing extensive and detailed threat reports. Effective reasoning enables the LLM to make informed recommendations, identify potential security risks, and suggest mitigation strategies based on the comprehensive analysis of the available data.
3.2. Data Sources and Tasks
@@
Employing a proprietary LLM like ChatGPT-4 (gpt, 2024b) is now a research standard in generating evaluation benchmarks in long-form responses (Yu et al., 2023; Kim et al., 2023; Zheng et al., 2024). We utilize the more recent OpenAI’s ChatGPT-4o (gpt, 2024c) in extracting our benchmark datasets, using suitable prompts. We also evaluate the quality of ChatGPT-4o output with human annotators to discard incorrect responses (annotation process and results discussed in Section 3.3). Below we explain the three different tasks of SECURE and various data sources:
Table 1. Sample MCQs and Answers from MAET and CWET
@@
3.2.1. Extraction Task
We frame the knowledge extraction task as a multiple-choice question answering (MCQ) task. The model is expected to answer questions without any given context, relying solely on its memory or training data. For this purpose, we focus on the MITRE ATT&CK (mit, 2024b) and CWE (Common Weakness Enumeration) (Project, 2024b) websites to create two datasets: MAET (Mitre Attack Extraction Task) and CWET (Common Weakness Extraction Task). Specifically, we utilize the attack patterns for ICS (mit, 2024c) and mitigation plans for ICS (mit, 2024a) to generate the questions. For CWE, we utilize the weaknesses belonging to the class CWE-1358 “CWE VIEW: Weaknesses in SEI ETF Categories of Security Vulnerabilities in ICS” (Project, 2024b), which enumerates all ICS-related security vulnerabilities.
We selected these resources due to their high relevance and quality in the domain of ICS cybersecurity. The MITRE ATT&CK framework is a globally recognized repository of tactics and techniques based on real-world observations, making it an authoritative source for constructing a benchmark dataset (mit, 2024b). Similarly, the CWE provides a community-developed list of software and hardware weaknesses that can become vulnerabilities, which is crucial for a comprehensive understanding of ICS security. The detailed and structured information from both MITRE ATT&CK and CWE ensures that the generated questions cover fundamental concepts and advanced technical details pertinent to ICS security. Below is an example prompt used in generating the questions for the MAET and CWET tasks in Prompt A (shortened). Using such prompts, we extract a total of 2036 questions. These questions test both basic and advanced understanding of ICS cybersecurity, ensuring the robustness of SECURE in evaluating the knowledge-intensive tasks by the LLM.

### v3

# SECURE: Benchmarking Large Language Models for Cybersecurity Advisory

@@ Abstract
Large Language Models (LLMs) have demonstrated potential in cybersecurity applications but have also caused lower confidence due to problems like hallucinations and a lack of truthfulness. Existing benchmarks provide general evaluations but do not sufficiently address the practical and applied aspects of LLM performance in cybersecurity-specific tasks. To address this gap, we introduce the SECURE (Security Extraction, Understanding & Reasoning Evaluation), a benchmark designed to assess LLMs performance in realistic cybersecurity scenarios. SECURE includes six datasets focussed on the Industrial Control System sector to evaluate knowledge extraction, understanding, and reasoning based on industry-standard sources. Our study evaluates seven state-of-the-art models on these tasks, providing insights into their strengths and weaknesses in cybersecurity contexts. We also offer recommendations for improving LLMs reliability as cyber advisory tools and release our benchmark datasets and framework for community use at https://github.com/aiforsec/SECURE.
Index Terms:
Large Language Models, Security, Cyber Advisory, Dataset.
I Introduction
@@
Recent breakthroughs in large language models (LLM) like OpenAI’s ChatGPT [1] have opened up their applications in many domains, including security [2]. These models, trained on vast datasets, can generate, understand, and reason across a multitude of domains [3]. Users can interact with these models in a conversational style through a simple web interface and obtain answers to their questions in a short time. Despite demonstrating high potential, LLMs are plagued by issues such as hallucinations [4] and truthfulness [5]. Hence, a standard evaluation benchmark is crucial to evaluate the reliability of such models. Several benchmarks like GLUE [6], MMLU [7], Helm [8] and KOLA [9], provide standard datasets and tasks to evaluate general-purpose understanding and capabilities of LLMs. GLUE assesses LLMs’ performance in understanding language, while MMLU and HELM offer a holistic evaluation across various domains like mathematics, history, computer science, and law. KOLA focuses on tasks designed to measure the cognitive abilities of LLMs.
Figure 1: Conversation between a user and a ChatGPT-3.5 on ICS security questions.
@@
Despite these advancements, there remains a significant gap in the evaluation of LLMs specifically tailored for security industries such as information security, network security, and critical infrastructure protection. Traditional benchmarks often fail to capture the practical and applied aspects of cybersecurity, leading to an incomplete assessment of LLM capabilities. These benchmarks typically focus on general language tasks and do not address specific challenges such as recognizing emerging threats, handling specialized terminology, or performing tasks like vulnerability assessment and incident response. More practical, domain-specific, and comprehensive evaluations are necessary to understand LLM performance in realistic cybersecurity scenarios [10]. In Figure 1, we demonstrate a conversation between ChatGPT-3.5 and a user based on our benchmark dataset (see Section III). ChatGPT-3.5 model was prompted to act as a security expert, and yet, all three responses, spanning different kinds of tasks, were incorrect, showing the unreliability of these models in cybersecurity. Especially, on the first task, we ask questions about vulnerabilities discovered in 2024. Even though the model is not trained on any such recent data, it confidently makes a decision, despite our instructions to return ’X’ when it is not confident.
Figure 2: Overview of the SECURE benchmark.
@@
To address this gap, we introduce a comprehensive benchmarking framework encompassing real-world cybersecurity scenarios, practical tasks, and applied knowledge assessments. We introduce Security ExtraCtion, Understanding & Reasoning Evaluation (SECURE) for large language models. This benchmark provides a holistic evaluation of LLMs performance in cybersecurity applications, ensuring they meet the high standards required for deployment in critical infrastructure environments, such as Industrial Control Systems (ICS). Figure 2 shows an overview of our proposed benchmark. SECURE incorporates knowledge modeling to design six different datasets (MAET: Mitre Attack Extraction Task, CWET: Common Weakness Extraction Task, KCV: Knowledge test on Common Vulnerabilities, VOOD: Vulnerability Out-of-Distribution task, RERT: Risk Evaluation Reasoning Task, and CPST: CVSS Problem Solving Task) focused on three types of knowledge evaluations: extraction, understanding, and reasoning. These datasets are sourced from standard sources, such as MITRE [11] [12], CVE [13], CWE [14], and Cybersecurity and Infrastructure Security Agency (CISA) [15].
@@
    Assessment of cybersecurity knowledge: We evaluate the cybersecurity knowledge of LLMs in assisting security analysts. To achieve this, we develop a series of benchmark tasks specifically for evaluating LLMs in the context of ICS cyber advisory (see Section III). These tasks are based on comprehensive protocol specifications detailing communication standards and procedures, historical cyber threats and attacks, CVEs with contextual information on exploitability and severity, and detailed remediation strategies, including patch notes and mitigation steps from MITRE.
  2. 2.
    Evaluation of LLMs: Recent works have shown that customized LLMs introduce a loss of safety measures [16], & increased hallucinations [17]. In addition, the computational & technical requirements of building customized LLMs can push researchers & industries to use commercial LLMs. However, before we can confidently claim the application of general-purpose LLMs in a specific domain, we first need to evaluate it on domain-specific tasks. For this, we first evaluate 7 open and closed source state-of-the-art models (ChatGPT-4 [1], ChatGPT-3.5 [18], Llama3-70b [19], Llama3-8b [20], Gemini-Pro [21], Mistral-7B [22] , Mixtral-8x7b [23]) on our benchmark tasks (see Section IV). We then customize Llama3-8b using Retrieval Augmented Generation (RAG) framework and fine-tuning and evaluate the LLMs on our benchmark (see Section VI).
  3. 3.
@@
    Insights and Recommendations: Our findings demonstrate that while LLMs demonstrate some capability in cybersecurity tasks, their use as advisory tools requires careful consideration. We provide insightful observations and propose recommendations to enhance their usability (see Section V).
  4. 4.
    Benchmark Dataset and Framework: We provide our benchmark datasets (MAET: Mitre Attack Extraction Task, CWET: Common Weakness Extraction Task, KCV: Knowledge test on Common Vulnerabilities, VOOD: Vulnerability Out-of-Distribution, RERT: Risk Evaluation Reasoning Task, and CPST: CVSS Problem Solving Task) and framework to the security community so that they can be openly used to evaluate future LLMs in various security applications.
II Background and Related Work
@@
Large Language Model (LLM) in Security: LLMs, trained on vast amounts of textual data, are capable of producing coherent and contextually relevant text. While earlier BERT-language models like SecureBERT [24] and CySecBERT [25] were used for language modeling in cybersecurity, the release of GPT (Generative Pre-trained Transformer) models has changed the nature of language models. There are two types of LLMs in the industry. Open Source LLMs like Llama [19] and Mixtral [22] make their models public so they can be fine-tuned for specific downstream tasks. Closed-source LLMs like ChatGPT [1] and Gemini [26] allow restricted access through APIs. The most notable language models in security are code-based LLMs such as CodeLlama [27], adapted to analyze and generate secure code. Open-sourced models have also been fine-tuned for several cybersecurity tasks like vulnerability detection [28, 29, 30, 31], program repair [32], IT operations [33], and security knowledge assistance [34].
@@
Evaluation Benchmark: While there are various benchmarks like GLUE [6], MMLU [7], Helm [8] and KOLA [9] for evaluating general-purpose LLMs, comprehensive cybersecurity-specific benchmarks remain limited. Existing approaches to evaluating LLMs in security tend to focus on factual knowledge rather than applied, practical cybersecurity tasks. For example: the CyberMetric dataset [35] comprises 10,000 questions sourced from cybersecurity standards. CyberBench [36] comprises ten datasets from different tasks namely, named-entity recognition, summarization, multiple choice, and classification. SecEval [37] evaluates cybersecurity knowledge in LLMs with 2000 multiple-choice questions across Software Security, Application Security, System Security, Web Security, Cryptography, Memory Safety, Network Security, and PenTest. SecQA [38] consists of multiple-choice questions based on the ”Computer Systems Security: Planning for Success” textbook to evaluate LLM’s understanding of security principles. NetEval [39] evaluates the knowledge of large language models in IT operational tasks within a multilingual context. OpsEval [40] also contains multi-choice questions designed for fault root cause analysis, operational script generation, and alert information summarization. Ullah et al. [41] tests LLMs on identifying and reasoning about software vulnerabilities using 228 code scenarios. Their findings suggest existing LLMs are unreliable in identifying vulnerabilities in source code. CTIBench [42] is a benchmark of four tasks to evaluate the ability of LLMs in cyber threat intelligence (CTI) landscape.
III Proposed Benchmark: SECURE
@@ III-A Modeling
We emphasize knowledge modeling in designing our evaluation benchmark. Knowledge, which includes both facts and skills, is a core indicator of intelligence [43]. Prior research shows that knowledge-intensive tasks can reliably evaluate the capabilities of LLMs [9, 44]. For cyber-advisory LLM, we aim to assess not only its ability to retrieve known facts but also gauge its proficiency in a) using context to answer questions and b) performing reasoning based on given some knowledge source. To create a robust evaluation benchmark, we focus on the following critical abilities of LLMs in handling knowledge-intensive tasks:
@@
Extraction: Knowledge extraction tasks are crucial for measuring the ability of a language model to access its vast knowledge base and accurately recall specific facts [44]. As a cybersecurity advisor, an LLM could be asked to retrieve information on various security architectures, incidents, best practices, and historical data on known vulnerabilities. Accurate extraction ensures the LLM can provide reliable and precise information quickly, which is essential for addressing security concerns, preventing breaches, and supporting security professionals in their decision-making processes.
@@
Understanding: While knowledge extraction tasks focus on recalling information, knowledge understanding tasks are designed to assess the cognitive abilities of a model [45]. This involves evaluating the model’s capability to discern the truthfulness of statements and comprehend underlying knowledge within a given context. For instance, an LLM might be tested on its ability to interpret the accuracy of security issues described in a report, thereby demonstrating its grasp of complex cybersecurity concepts and scenarios. Effective understanding ensures that the LLM can accurately interpret and respond to nuanced security challenges.
@@
Reasoning: Knowledge reasoning tasks aim to evaluate the problem-solving capabilities of LLMs [46]. This is particularly important in the cyber-advisory role of LLMs, as they need to assist security professionals in reading, analyzing, and summarizing extensive and detailed threat reports. Effective reasoning enables the LLM to make informed recommendations, identify potential security risks, and suggest mitigation strategies based on the comprehensive analysis of the available data.
III-B Data Sources and Tasks
@@
Employing a proprietary LLM like ChatGPT-4 [1] is now a research standard in generating evaluation benchmarks in long-form responses [9, 47, 48]. We utilize the more recent OpenAI’s ChatGPT-4o [49] in extracting our benchmark datasets, using suitable prompts. We also evaluate the quality of ChatGPT-4o output with human annotators to discard incorrect responses (annotation process and results discussed in Section III-C). Below we explain the three different tasks of SECURE and various data sources:
TABLE I: Sample MCQs and Answers from MAET and CWET
@@
III-B1 Extraction Task
We frame the knowledge extraction task as a multiple-choice question answering (MCQ) task. The model is expected to answer questions without any given context, relying solely on its memory or training data. For this purpose, we focus on the MITRE ATT&CK [50] and CWE (Common Weakness Enumeration) [14] websites to create two datasets: MAET (Mitre Attack Extraction Task) and CWET (Common Weakness Extraction Task). Specifically, we utilize the attack patterns for ICS [11] and mitigation plans for ICS [12] to generate the questions. For CWE, we utilize the weaknesses belonging to the class CWE-1358 “CWE VIEW: Weaknesses in SEI ETF Categories of Security Vulnerabilities in ICS” [14], which enumerates all ICS-related security vulnerabilities.
We selected these resources due to their high relevance and quality in the domain of ICS cybersecurity. The MITRE ATT&CK framework is a globally recognized repository of tactics and techniques based on real-world observations, making it an authoritative source for constructing a benchmark dataset [50]. Similarly, the CWE provides a community-developed list of software and hardware weaknesses that can become vulnerabilities, which is crucial for a comprehensive understanding of ICS security. The detailed and structured information from both MITRE ATT&CK and CWE ensures that the generated questions cover fundamental concepts and advanced technical details pertinent to ICS security. Below is an example prompt used in generating the questions for the MAET and CWET tasks in Prompt A (shortened). Using such prompts, we extract a total of 2036 questions. These questions test both basic and advanced understanding of ICS cybersecurity, ensuring the robustness of SECURE in evaluating the knowledge-intensive tasks by the LLM.